# EmotionIC: emotional inertia and contagion-driven dependency modeling for emotion recognition in conversation

### 25 Dec 2023 v4

@@ Abstract
Emotion Recognition in Conversation (ERC) has attracted growing attention in recent years as a result of the advancement and implementation of human-computer interface technologies. In this paper, we propose a novel approach to dependency modeling driven by Emotional Inertia and Contagion (EmotionIC) for ERC task. Our EmotionIC consists of three main components, i.e., Identity Masked Multi-Head Attention (IMMHA), Dialogue-based Gated Recurrent Unit (DiaGRU), and Skip-chain Conditional Random Field (SkipCRF). Compared to previous ERC models, EmotionIC can model a conversation more thoroughly at both the feature-extraction and classification levels. The proposed model attempts to integrate the advantages of attention- and recurrence-based methods at the feature-extraction level. Specifically, IMMHA is applied to capture identity-based global contextual dependencies, while DiaGRU is utilized to extract speaker- and temporal-aware local contextual information. At the classification level, SkipCRF can explicitly mine complex emotional flows from higher-order neighboring utterances in the conversation. Experimental results show that our method can significantly outperform the state-of-the-art models on four benchmark datasets. The ablation studies confirm that our modules can effectively model emotional inertia and contagion.
@@ 1 Introduction
Figure 1: An example of contextual dependency modeling at both the feature-extraction and classification levels. The dashed and solid line represent the intra- and inter-speaker information transmissions, respectively.
In summary, we propose a novel approach to dependency modeling driven by Emotional Inertia and Contagion (EmotionIC) in this paper, which can adequately model contexts from the feature-extraction and classification levels. Figure 1 is an example of modeling context dependency at both the feature-extraction and classification levels. At the feature-extraction level, we design an Identity Masked MHA (IMMHA) to capture intra- and inter-speaker dependencies in the global contexts from two features subspaces, respectively; to further refine the contextual dependencies, we devise a speaker- and position-aware Dialogue GRU (DiaGRU). DiaGRU introduces the emotional tendencies of the current speaker and interlocutor at the previous moment into the single GRU cell. At the classification level, by introducing skip connections in CRF, we craft a novel structure called Skip-chain CRF (SkipCRF) to explicitly capture the emotional flows in the conversation. SkipCRF takes into account higher-order contextual dependencies from intra- and inter-speaker, and can model the complex emotional interactions of different participants in the conversation. It is worth noting that instead of using additional softmax layer, we directly utilize CRF for final emotion classification. Experimental results demonstrate the superiority of our model compared with state-of-the-art models, and several studies are conducted to illustrate the effectiveness of each module of EmotionIC. Put it briefly, Our main contributions are summarized as follows:
@@ 3 Our Approach
Figure 2: The architecture of our EmotionIC. Firstly, the global and local context dependencies are extracted through IMMHA and DiaGRU, respectively. Then, we concatenate global context features, local context features and utterance features. Note that combining utterance features that are not processed by IMMHA and DiaGRU is to prevent the original semantics of utterances with weak context dependencies from being obscured. Finally, the emotional flows in the conversation are captured at the classification level through SkipCRF to obtain the optimal emotion sequence.
@@ 3.1 Preliminaries
Utterance Moment Function. To distinguish contextual dependencies from different participants, two utterance moment functions, i.e., $\bm{s}(\cdot)$ and $\bm{o}(\cdot)$, are defined in conversation. Here, $\bm{s}(\cdot)$ is utilized to output the moment of the previous utterance that belongs to the same speaker as the current utterance and is the nearest to that utterance; $\bm{o}(\cdot)$ is employed to output the moment of the previous utterance that is spoken by the interlocutor and is the nearest to the current utterance. For instance, assuming the existence of the current utterance $u_{t}^{p_{i}}$, then $u_{\bm{s}(t)}^{p_{i}}$ represents the nearest and previous utterance uttered by the speaker $p_{i}$, and $u_{\bm{o}(t)}^{p_{j}}$ indicates the most recent utterance spoken by the interlocutor $p_{j}$.
Figure 3: The illustration of utterance moment function. $u_{t-2}^{p_{i}}$ is the nearest and previous utterance uttered by the speaker $p_{i}$, so it can also be denoted by $u_{\bm{s}(t)}^{p_{i}}$. Similarly, $u_{t-1}^{p_{j}}$ can be also denoted by $u_{\bm{o}(t)}^{p_{j}}$.
@@ 3.2 Identity Masked Multi-head Attention
Figure 4: The network structure of IMMHA. $M_{s}$ and $M_{o}$ are two mask matrices that mask contextual dependencies from other participants and the current participant, respectively, as well as future information. $\mathbf{MatMul}$ indicates the matrix multiplication operation, and $\mathbf{Mask}$ is the masking operation.
@@ 3.4 Skip-chain Conditional Random Field
(a) Before Elimination
(b) After Elimination
Figure 6: An example of CRF modeling for a three-person conversation scenario. The nodes with different colors represent different participants’ node features. The dotted and solid lines represent the weights of self- and other-dependent features, respectively. The filled nodes represent the ground-truth labels, and the red line connects the optimal emotion sequence and represents the optimal path.
@@ 4.1 Evaluation Metrics and Datasets
Table 1: The statistics for these four emotion datasets. #Conversation and #Utterance denote the number of conversations and utterances, respectively.
@@ 4.2 Implementation Details
Table 2: Partial hyperparameter settings for distinct datasets. LR-ext and LR-cls denote the learning rates for the feature-extraction and classification levels, respectively. BS and DR is the batch size and dropout rate, respectively. D-IMMHA and D-DiaGRU indicate the network depths for IMMHA and DiaGRU modules, respectively.
@@ 5.1 Comparison With Baseline Methods
Table 3: The experimental results of EmotionIC and baseline models on four datasets. All baseline results are from the original paper and only textual modality is used.
@@
    MELD. The utterances in the MELD dataset are much shorter than those in the IEMOCAP dataset, which means that emotional modeling is highly context-dependent. In addition, the fact that there are often two or more speakers in a conversation and that the amount of statements uttered by each participant is relatively small makes it difficult to model based on emotional inertia and contagion. While our EmotionIC reach the best F1 score, even surpassing that of COSMIC which adds additional commonsense knowledge. We attribute the improvement to our contextual modeling at the classification level, i.e., SkipCRF.
@@ 5.2 Analysis for Confusion Matrices
(a) IEMOCAP Dataset
(b) DailyDialog Dataset
(c) MELD Dataset
(d) EmoryNLP Dataset
Figure 7: Confusion matrices of the testing set on four benchmark datasets. The horizontal and vertical coordinate indicate the ground-truth and predicted emotion, respectively. Note that the majority class Neutral in the DailyDialog dataset is removed.
@@ 5.3 Result for Each Emotion
(a) IEMOCAP Dataset
(b) MELD Dataset
Figure 8: F1 score for each emotion on the MELD and IEMOCAP datasets. Note that Hap represents for the first three-letter abbreviation for Happy, and so on for other emotions.
@@ 5.4 Ablation Studies
Table 4: The impact of removing different modules. - indicates removal of the corresponding module.
@@
Table 5: The impact of replacing different modules. - indicates removal of the corresponding module, + indicates replacement with another module.
@@ 5.5 Effectiveness of SkipCRF
Figure 9: A case study on the MELD dataset. Prediction (SkipCRF) denotes the use of SkipCRF for emotion classification, and Prediction (Softmax) denotes the replacement of SkipCRF with the Softmax layer for emotion classification.
@@ 5.6 Class Imbalance in Both the Dialogue and Dataset
Figure 10: The effect of the class imbalance in the dialogue and that in the dataset on the performance. Note that we explore on the DailyDialog (- Neutral) dataset since it has an extreme class imbalance problem.
@@ 5.7 Comparison of Modeling Emotional Inertia and Contagion
Table 6: F1 scores of EmotionIC on the samples with emotional inertia and contagion. We adopt micro F1 score to record the results on the DailyDialog dataset and adopt weighted F1 score on the other datasets.
@@
Table 7: F1-score of EmotionIC on the samples with emotional inertia and contagion for different emotions. These results are derived from the DailyDialog (- Neutral) dataset.

### 20 Mar 2024 v5

# Received 23 March 2023/Revised 28 August 2023/Revised 16 October 2023/Accepted 10 November 2023

@@ Abstract
Emotion Recognition in Conversation (ERC) has attracted growing attention in recent years as a result of the advancement and implementation of human-computer interface technologies. In this paper, we propose an emotional inertia and contagion-driven dependency modeling approach (EmotionIC) for ERC task. Our EmotionIC consists of three main components, i.e., Identity Masked Multi-Head Attention (IMMHA), Dialogue-based Gated Recurrent Unit (DiaGRU), and Skip-chain Conditional Random Field (SkipCRF). Compared to previous ERC models, EmotionIC can model a conversation more thoroughly at both the feature-extraction and classification levels. The proposed model attempts to integrate the advantages of attention- and recurrence-based methods at the feature-extraction level. Specifically, IMMHA is applied to capture identity-based global contextual dependencies, while DiaGRU is utilized to extract speaker- and temporal-aware local contextual information. At the classification level, SkipCRF can explicitly mine complex emotional flows from higher-order neighboring utterances in the conversation. Experimental results show that our method can significantly outperform the state-of-the-art models on four benchmark datasets. The ablation studies confirm that our modules can effectively model emotional inertia and contagion.
@@ 1 Introduction
Figure 1: Example of contextual dependency modeling at both the feature-extraction and classification levels. The dashed and solid line represent the intra- and inter-speaker information transmissions, respectively.
In summary, we propose an Emotional Inertia and Contagion-driven dependency modeling approach (EmotionIC) in this paper, which can adequately model contexts from the feature-extraction and classification levels. Figure 1 is an example of modeling context dependency at both the feature-extraction and classification levels. At the feature-extraction level, we design an Identity Masked MHA (IMMHA) to capture intra- and inter-speaker dependencies in the global contexts from two features subspaces, respectively; to further refine the contextual dependencies, we devise a speaker- and position-aware Dialogue GRU (DiaGRU). DiaGRU introduces the emotional tendencies of the current speaker and interlocutor at the previous moment into the single GRU cell. At the classification level, by introducing skip connections in CRF, we craft a novel structure called Skip-chain CRF (SkipCRF) to explicitly capture the emotional flows in the conversation. SkipCRF takes into account higher-order contextual dependencies from intra- and inter-speaker, and can model the complex emotional interactions of different participants in the conversation. It is worth noting that instead of using additional softmax layer, we directly utilize CRF for final emotion classification. Experimental results demonstrate the superiority of our model compared with state-of-the-art models, and several studies are conducted to illustrate the effectiveness of each module of EmotionIC. To put it briefly, our main contributions are summarized as follows:
@@ 3 Our approach
Figure 2: Architecture of our EmotionIC. Firstly, the global and local context dependencies are extracted through IMMHA and DiaGRU, respectively. Then, we concatenate global context features, local context features and utterance features. Note that combining utterance features that are not processed by IMMHA and DiaGRU is to prevent the original semantics of utterances with weak context dependencies from being obscured. Finally, the emotional flows in the conversation are captured at the classification level through SkipCRF to obtain the optimal emotion sequence.
@@ 3.1 Preliminaries
Utterance moment function. To distinguish contextual dependencies from different participants, two utterance moment functions, i.e., $\bm{s}(\cdot)$ and $\bm{o}(\cdot)$, are defined in conversation. Here, $\bm{s}(\cdot)$ is utilized to output the moment of the previous utterance that belongs to the same speaker as the current utterance and is the nearest to that utterance; $\bm{o}(\cdot)$ is employed to output the moment of the previous utterance that is spoken by the interlocutor and is the nearest to the current utterance. As shown in Figure 3, assuming the existence of the current utterance $u_{t}^{p_{i}}$, then $u_{\bm{s}(t)}^{p_{i}}$ represents the nearest and previous utterance uttered by the speaker $p_{i}$, and $u_{\bm{o}(t)}^{p_{j}}$ indicates the most recent utterance spoken by the interlocutor $p_{j}$.
Figure 3: Illustration of utterance moment function. $u_{t-2}^{p_{i}}$ is the nearest and previous utterance uttered by the speaker $p_{i}$, so it can also be denoted by $u_{\bm{s}(t)}^{p_{i}}$. Similarly, $u_{t-1}^{p_{j}}$ can be also denoted by $u_{\bm{o}(t)}^{p_{j}}$.
@@ 3.2 Identity masked multi-head attention
Figure 4: Network structure of IMMHA. $M_{s}$ and $M_{o}$ are two mask matrices that mask contextual dependencies from other participants and the current participant, respectively, as well as future information. $\mathbf{MatMul}$ indicates the matrix multiplication operation, and $\mathbf{Mask}$ is the masking operation.
@@ 3.4 Skip-chain conditional random field
Figure 6: Example of CRF modeling for a three-person conversation scenario (a) before and (b) after eliminating skip-chain connections. The nodes with different colors represent different participants’ node features. The dotted and solid lines represent the weights of self- and other-dependent features, respectively. The filled nodes represent the ground-truth labels, and the red line connects the optimal emotion sequence and represents the optimal path.
@@ 4.1 Evaluation metrics and datasets
Table 1: Statistics for these four emotion datasets. #Conversation and #Utterance denote the number of conversations and utterances, respectively.
@@ 4.2 Implementation details
Table 2: Partial hyperparameter settings for distinct datasets. LR-ext and LR-cls denote the learning rates for the feature-extraction and classification levels, respectively. BS and DR are the batch size and dropout rate, respectively. D-IMMHA and D-DiaGRU indicate the network depths for IMMHA and DiaGRU modules, respectively.
@@ 5.1 Comparison with baseline methods
Table 3: Experimental results of EmotionIC and baseline models on four datasets. All baseline results are from the original paper and only textual modality is used.
@@
    MELD. The utterances in the MELD dataset are much shorter than those in the IEMOCAP dataset, which means that emotional modeling is highly context-dependent. In addition, the fact that there are often two or more speakers in a conversation and that the amount of statements uttered by each participant is relatively small makes it difficult to model based on emotional inertia and contagion. Our EmotionIC reaches the best F1 score, even surpassing that of COSMIC which adds additional commonsense knowledge. We attribute the improvement to our contextual modeling at the classification level, i.e., SkipCRF.
@@ 5.2 Analysis for confusion matrices
Figure 7: Confusion matrices of the testing set on the (a) IEMOCAP, (b) DailyDialog, (c) MELD, and (d) EmoryNLP datasets. The horizontal and vertical coordinate indicate the ground-truth and predicted emotion, respectively. Note that the majority class Neutral in the DailyDialog dataset is removed.
@@ 5.3 Result for each emotion
Figure 8: F1 score for each emotion on the (a) IEMOCAP and (b) MELD datasets. Note that Hap represents for the first three-letter abbreviation for Happy, and so on for other emotions.
@@ 5.4 Ablation studies
Table 4: Impact of removing different modules. - indicates removal of the corresponding module.
@@
Table 5: Impact of replacing different modules. - indicates removal of the corresponding module, + indicates replacement with another module.
@@ 5.5 Effectiveness of SkipCRF
Figure 9: Case study on the MELD dataset. Prediction (SkipCRF) denotes the use of SkipCRF for emotion classification, and Prediction (Softmax) denotes the replacement of SkipCRF with the Softmax layer for emotion classification.
@@ 5.6 Class imbalance in both the dialogue and dataset
Figure 10: Effect of the class imbalance in the dialogue and that in the dataset on the performance. Note that we explore on the DailyDialog (- Neutral) dataset since it has an extreme class imbalance problem.
@@ 5.7 Comparison of modeling emotional inertia and contagion
Table 6: F1 scores of EmotionIC on the samples with emotional inertia and contagion for different datasets. We adopt micro F1 score to record the results on the DailyDialog dataset and adopt weighted F1 score on the other datasets.
@@
Table 7: F1 scores of EmotionIC on the samples with emotional inertia and contagion for different emotions. These results are derived from the DailyDialog (- Neutral) dataset.